OpenStudy (anonymous):

if you were to fall into a 100m deep crater (on the moon) how fast would you be going when you hit, and how long would it take you.--What are the steps in figuring this out?

OpenStudy (amistre64):

the steps are in figuring out with the gravitational constant is for the moon

OpenStudy (amistre64):

or rather, the constant acceleration due to gravity on the moon

OpenStudy (amistre64):

Unless youre given a formula; then you would have to calculate the mass of the moon. its radius and such

OpenStudy (amistre64):

google, the terms and see if you can find the formula for the moon

OpenStudy (amistre64):

it should be like an 8th or so but i cant recall

OpenStudy (anonymous):

if the moons radius is 1738000 and the mass is 7.35e22 and the acceleration due to gravity is 1.63

OpenStudy (amistre64):

acceleration due to gravity is the kicker

OpenStudy (amistre64):

is that in meters of feet?

OpenStudy (amistre64):

1.63 is meters

OpenStudy (amistre64):

so its the right units

OpenStudy (amistre64):

-1.63 is constant; meaning that your being pulled down at a constant rate a(t) = -1.63 v(t) = -1.63t + C ; which at t=0, C=0 s(t) = -.815 t^2 + Vi + Hi = position then

OpenStudy (amistre64):

lets assume you fall instead of are pushed; Vi = 0 and it says you are 100 for you initial height Hi

OpenStudy (amistre64):

s(t) = -.815 t^2 + 100; when this =0 you are on the ground -.815 t^2 + 100 = 0 when t is about 11.1 seconds

OpenStudy (amistre64):

How fast is v(t) v = -1.63 (11.1) = about 18.1 meters/sec

OpenStudy (amistre64):

in the downward direction; so speed is an absolute value